Suppose the following data represents the produced units by applied new production systems, sold price for each unit in each system, the required labor hours for production, the labor hour cost, and the overhead for each production process. Products AED/unit Old System production 4000 100 New system production 6000 120 Old System Labor hours 400 10 New System Labor hours 500 12 Old System overheads 1000 New System overheads 1500 a) Calculate labor productivity for the two systems and Calculate the labor cost productivity for the two systems b) Calculate the multifactor productivity for the two systems c) Compare between the productivity in each measure and comment.
Answer #1

Labor Productivity:

                For old system:

                No of products = 4000

                Labor hours required = 400

Hence Labor productivity = (Labor hours/no of products) = 4000/400 = 10 units per hour

For new system:

No of products = 6000

Labor hours required = 500

Hence Labor productivity = (Labor hours/no of products) = 6000/500 = 12 units per hour

Labor cost productivity

                For old system:

                No of products = 4000

                Labor hours required = 400

                Labor cost per hour = $10

                Total labor cost = 400*10 = 4000

                Hence Labor cost productivity = labor cost/no of products = 4000/4000 = $1 per unit

                For new system:

                No of products = 6000

                Labor hours required = 500

                Labor cost per hour = $12

                Total labor cost = 500*12 = 6000

                Hence Labor cost productivity = labor cost/no of products = 6000/6000 = $1 per unit

Multifactor Productivity:

Multifactor productivity = Units of Output / (Units of Labor + Units of Capital + Units of overhead)

                For old system:

                No of products = 4000

                Labor hours required = 400

                Labor cost per hour = $10

                Total labor cost = 400*10 = 4000

                Overhead cost = $1000

                Cost per unit = $100

                For 4000 units cost = $400000

                Multifactor productivity = 4000/($4000+$400000+$1000) = 4000/(405000) units per dollar

                If we count in dollar per unit: it will be 405000/4000 = $101.25 per unit.

                For old system:

                No of products = 6000

                Labor hours required = 500

                Labor cost per hour = $12

                Total labor cost = 500*12 = 6000

                Overhead cost = $1500

                Cost per unit = $120

                For 6000 units cost = $720000

                Multifactor productivity = 6000/($6000+$720000+$1500) = 6000/(727500) units per dollar

                If we count in dollar per unit: it will be 727500/6000 = $121.25 per unit.

IF YOU LIKE THE ANSWER, PLEASE GIVE AN UP-VOTE OR THUMB UP. THIS WILL ENCOURAGE ME TO ANSWER MORE!!